Abundance 108 CH200C 164 151 206 234 262 276290 304 50 347 220 248210304 178 347 192 | 332 With 60 80 100 120 140 160 180 200 Show the fragmentation pattern.
Answer #1

nov 10 1 14 thin VDAY m/2=206 13 Els mea 118 Got m/2 = 184 m/2=151 mrm/2=108 MY M12-92 T111 164 it lose (14) each time that m